Marvin is currently three times as old as Al. In twelve years, Marvin will be six less than twice Al’s age. How old is Marvin today?
Answer by ikleyn(52794)

.

Al  = x years old.

Marvin = 3x years old.


In 12 years

    3x + 12 = 2*(x+12) - 6.


It is your basic equation from the condition.


Simplify and solve


    3x + 12 = 2x + 24 - 6

    3x + 12 = 2x + 18

    3x - 2x = 18 - 12

       x    = 6.


So, Al is 6 years old now;  Marwin is 3*6 = 18 years old.

Solved.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.